A chemical distributor can purchase raw material from Mexico at three different price policies. Complete the shaded cells and answer the following questions.

Policy Discount Quantity (lbs) Discount (%) Discount Price (C) per lb EOQ Feasible/Not feasible Optimal Order Size for the discount range policy
1 0 to 199 No discount 20 565 ______________ _____________
2 200 to 800 5 19 580 Feasible 580
3 801 and above 10 18 ___ ______________ ___________

(Annual demand = 20000 lbs; Ordering cost (setup) = $40; Holding cost = 25% of cost)

a) What is the optimal order size for discount policies 1 and 2?
b) What is the total cost of the overall best policy for this chemical distributor?
Answer #1

Given Data,

For Supplier A,

[ Ch] Holding Cost = 25% of the Unit Cost per unit per Year,

[Co ]Ordering Cost = USD 40 Per Order,

[ D ]Annual Demand = 20,000 lbs

For Policy – 1 – Not Feasible

Formula for Optimal Order Quantity =

Square Root of [( 2 X D X Co)/ Ch],

Or, Optimal Order Quantity = Square Root of [ (
2 X 20,000 X 40 )/ (0.25 X 20)]

Or, Optimal Order Quantity = Square Root of [ 320000] = 565.685 = 566 lbs

Item cost could vary as per the table data, Considering each Item cost, if we calculate Optimal order quantity, then the result will fall in between 200 to 800, thus let me consider, for any order in between 200 to 800, Price will come USD19 per lbs,

Same way Policy – 3 also Not Feasible

For Policy – 2 – Best Policy

Formula for Optimal Order Quantity =

Square Root of [( 2 X D X Co)/ Ch],

Or, Optimal Order Quantity = Square Root of [ (
2 X 20,000 X 40 )/ (0.25 X 19)]

Or, Optimal Order Quantity = Square Root of [ 336842] = 580.38 = 580 lbs

So, for the Optimal order quantity 580 Lbs, Firm should pay the price per unit = USD 19 per lbs,

Number of Order = Annual Demand / Optimal Order Quantity

Or, Number of Order = 20,000 / 580 = 34.48

Total Annual Ordering Cost = Number of Order X Ordering Cost per Order = 34.48 X USD 40 = USD 1379.31

Total Annual Holding Cost = Average Units X Holding Cost per unit per Year

Or, Total Annual Holding Cost = (580/2) X USD (0.25 X 19 ) = USD 1377.5,

Total Annual Purchasing Cost = Order Quantity X Number of Order X Purchase cost per Unit = 580 X 34.48 X USD 19 = USD 380000

Total Annual Cost = Total Annual Ordering Cost + Total Annual Holding Cost + Total Annual Purchasing Cost

Or, Total Annual Cost at the Optimal Behavior = USD 1379.31+ USD 1377.5 + USD 380000 = USD 382756.80